# Refined Policy Distillation: From VLA Generalists to RL Experts

### v1

@@ IV-A RPD Variants and Baselines
We compare three different variants of RPD on ManiSkill2’s Cube task:
@@
All RPD variants are trained with PPO and constant hyperparameters which were adopted without modification from the ManiSkill2 baseline [31]. The only exception is the batch size that had to be adjusted to fit the VLA on the same GPU as the RL training loop. Due to compute constraints we distilled from OpenVLA only with RPD-MSE.
We train policies with vanilla PPO and PPD as baselines and use the PPO implementation from CleanRL [32] that is included in ManiSkill2. As there is no source code publicly available for PPD, we use our own implementation, which extends CleanRL’s PPO. To compute the KL divergence PPD requires for its distillation loss, we sample ten actions from the VLA and fit a multivariate Gaussian distribution to it. All hyperparameters are set to ManiSkill2’s defaults. The weighting factor of PPD is set to $\lambda=1.0$.
Fig. 3 depicts the validation success rates during training for all RPD loss variants and baselines mentioned above. For reference, it also includes the performance of the VLAs after fine-tuning on the ManiSkill2 dataset previously presented in Tab. I.
The evaluation shows that RPD-MSE outperforms all other variants, closely followed by RPD-L1. The distilled policy quickly surpasses Octo and OpenVLA to become a refined policy. RPD-MSE with OpenVLA performs slightly worse than the distillation with Octo. The PPO baseline also converges to similar success rates as RPD-MSE and RPD-L1 of around 80% but learns substantially slower and shows much larger fluctuations in the training process. RPD-BC slightly surpasses Octo’s performance threshold but stays at that level and does not refine the policy. Lastly, PPD fails to lift the policy above the Octo baseline.
The reason why our PPD baseline does not manage to learn the task successfully is likely that it requires the teacher and student policy distributions to be of similar shape for the KL divergence to converge. However, we found that the action distributions of the VLA policies are sometimes bimodal, which results in non-optimal parameters for the approximated normal distributions that we use to compute the KL divergence in PPD. We also speculate that in our setup PPD depends strongly on the teacher’s performance. Lastly, PPD might also be sensitive to PPO hyperparameters, which we have not investigated. We did, however, change the parameter $\lambda$, which did not lead to substantially different results.
@@
Figure 3: Average success rates during training for vanilla PPO, PPD, and three different RPD variants: RPD-MSE, RPD-L1, and RPD-BC, which all distill from Octo over five runs with different seeds. indicate standard deviations. We also evaluated RPD-MSE on OpenVLA. Due to computing constraints, we could only perform a single training run. The performance of the fine-tuned VLAs is indicated by dashed lines.
@@ IV-B RPD on Different Tasks
Figure 4: Success and reward curves of RPD with Octo and vanilla PPO for both dense and sparse rewards on the six different ManiSkill2 tasks tha part of the fine-tuning dataset. The values are averaged over five runs with different seeds and are recorded in an evaluation environment. Standard deviations are indicated by shaded areas. All training runs were performed with the same hyperparameters taken from the ManiSkill2 PPO baseline. For all tasks, RPD outperforms Octo quickly and converges faster than vanilla PPO. In some cases, it even finds good policies when vanilla PPO fails. This effect increases for sparse rewards.
Fig. 4 shows the training curves for RPD-MSE with Octo on six different manipulation tasks that were part of the training dataset. In all six cases, RPD learns faster than the PPO baseline and often converges to higher success rates. This increased performance can also be seen in the reward plots. For RollBall and StackCube, the ManiSkill2 PPO hyperparameters fail to solve the tasks whereas RPD converges at least to partial success using the very same parameters. The decreased standard deviation also shows that RPD is generally more robust than the vanilla PPO baseline. These results demonstrate that RPD improves PPO’s training speed consistently and over many tasks while surpassing its final performance. It also refines the distilled policy on dense rewards.
Figure 5: Success and reward curves for RPD applied to Octo and vanilla PPO on tasks that are not part of Octo’s fine-tuning dataset. The curves show average results for five training runs on different seeds and are recorded in evaluation environments. The shaded areas indicate standard deviations. Note that the Octo baseline does not represent the correct reward in the dense reward setting as it solves the tasks without tool use. See sec. III for details.
@@ IV-C Sparse Rewards and Sample Efficiency
ManiSkill2 defines its sparse rewards as follows: $-1$ for failed cases, $1$ for success, and 00 otherwise. This means that non-zero reward is only given at the end of the episode, which is different from dense rewards where every action usually receives a non-zero reward that is shaped to guide the agent towards a certain behavior. Sparse reward problems are often a big challenge for RL algorithms, as relevant information about the agent’s performance is only obtained once the task has been solved successfully at least one time. Thus, in the beginning, the agent needs to complete an episode successfully just through random exploration, which makes the training process very inefficient compared to RL with dense rewards. As a result, it can occur that an agent learns a task perfectly fine with dense rewards but fails in the corresponding sparse reward setting, as can be seen for PPO in Fig. 4 for some of the tasks.
Fig. 4 also shows training runs with the six selected ManiSkill2 tasks for sparse rewards. Note that the resulting reward values cannot be directly compared to the dense reward setting. The only modification we applied to the hyperparameters from the ManiSkill2 baselines was to set $\gamma$ from $0.8$ to $0.99$, which avoids vanishing returns for early episode steps. In all six cases, RPD on sparse rewards consistently outperforms vanilla PPO. In most tasks, the baseline completely fails to learn the task, whereas RPD succeeds. It makes sense that RPD can realize its full potential on sparse rewards as it is even more important in this setting to have a teacher policy that enables the student to successfully complete episodes and, thus, retrieve environment rewards that drive the learning process. In summary, RPD drastically improves sample efficiency for on-policy RL in both sparse and dense reward cases.
@@ IV-D Generalization to New Environments
We also evaluated RPD with Octo on tasks that were held out of the VLA fine-tuning dataset to test its generalization capabilities. Ultimately, RPD is limited by the performance of the teacher and, in this case, by the cross-task generalization of the VLA. Still, it is worth investigating to what extent the degraded VLA performance will influence the RPD training. We study this question by looking at task variations, which means they are out-of-distribution with respect to the fine-tuning dataset. More specifically, we evaluate RPD’s performance on the hold-out tasks PullCubeTool and PokeCube as well as on a variation of the PushCube task with a different camera perspective.
@@
VLAs have traditionally been very sensitive to camera adjustments, leading to severe performance degradation [9]. To counteract performance drops VLAs often require fine-tuning with camera data from the new angle, camera model, etc. This problem also occurs with our fine-tuned versions of Octo and OpenVLA: When we only slightly change the camera angle in the PushCube task, as shown in Fig. 6, the success rates of OpenVLA drops from 27% to 4.5%. Octo even drops from 67% to 0%. The latter is likely a result of the less powerful vision encoder that is utilized in Octo which may overfit the training data.
Fig. 7 shows the success rates of vanilla PPO and RPD-MSE for both Octo and OpenVLA on the changed camera perspective from Fig. 6 on the PushCube task. PPO performs similar on our task variation as in the original PushCube task with sparse rewards, see Fig. 4. This is expected because the agent’s visual encoder is learned from scratch in the RL training. Accordingly, there is no bias towards a specific camera angle as long as the provided information is sufficient to solve the task. Both RPD variants manage to outperform PPO.
@@
Figure 7: Average evaluation success rates of vanilla PPO and RPD-MSE with both Octo and OpenVLA on a variation of the PushCube task with sparse rewards over five different seeds. Shaded areas indicate standard deviations. The camera perspective was changed as shown in Fig. 6 to investigate the distillation performance of the degraded VLAs.
@@ IV-E Limitations
The main limitation of RDP is the performance of the VLA on the task at hand. If the VLA cannot provide any meaningful actions due to domain shift or because the task was not included in the training dataset, RPD can be seen as a new exploration component in the loop. Depending on the resulting action distribution, the VLA may provide bias that is useful for learning the task or could improve the exploration of the environment.
However, if the actions generated by the VLA cannot provide such helpful bias, RPD may also decrease the RL agent’s performance. As reported for the training with dense rewards on the hold-out tasks, if the task and no variation of it have ever been seen by the VLA, its actions hinder the PPO training in RPD, and vanilla PPO performs better. hen PPO is already tuned to a task or the task is very easy, PPO is already sample-efficient and RPD may only provide limited improvement.
@@ V Conclusion and Future Work
This work introduced RPD, a novel on-policy RL-based approach for distilling generalist VLAs into refined expert policies. To the best of our knowledge, our work is the first that distills VLAs into compact expert policies. Our experiments show that RPD not only makes the RL training more efficient by guiding its exploration through the actions of the VLA but also that it can leverage task knowledge from the VLA even in situations where the generalist policy fails, such as when the camera perspective changes. Thus, RPD can help to improve the usability and accessibility of VLAs.
Despite the reported efficiency gains in the RL process, RPD still requires training in simulation, which is why we fine-tuned two state-of-the-art generalist VLA policies, Octo and OpenVLA, on simulation datasets. As a side contribution, our work shows that VLAs can be successfully utilized in experiments in simulated environments. This provides a promising perspective for scientists who have no access to a physical robot setup to conduct research on VLAs. An interesting direction for future research is to combine RPD with sim-to-real methods [33] for deployment on real robots.
@@ Acknowledgement
This work is supported by the project GeniusRobot funded by the German Federal Ministry of Education and Research (BMBF grant no. 01IS24083). The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the BayernKI project no. v108be. BayernKI funding is provided by Bavarian state authorities.

### v2

@@ IV-A RPD Variants and Baselines
We compare three different variants of RPD on ManiSkill3’s PickCube task:
@@
We use the same policy architecture for all tasks, RPD variants, and baselines. All RPD variants are trained with PPO and constant hyperparameters that were adopted without modification from the ManiSkill3 baseline [31]. The only exception is the batch size, which had to be adjusted to fit the VLA on the same GPU as the RL training loop. Due to computational constraints, we distilled from OpenVLA only with RPD-MSE.
We train policies with vanilla PPO and PPD [23] as baselines and use the PPO implementation from CleanRL [32] that is included in ManiSkill3. As there is no source code publicly available for PPD, we use our own implementation, which extends CleanRL’s PPO. To compute the KL divergence required by PPD for its distillation loss, we sample ten actions from the VLA and fit a multivariate Gaussian distribution to it. All hyperparameters are set to ManiSkill3’s defaults. The weighting factor of PPD is set to $\lambda=1.0$.
Fig. 3 depicts the validation success rates during training for all RPD loss variants and baselines mentioned above. For reference, it also includes the performance of the VLAs after fine-tuning on the ManiSkill3 dataset presented in Tab. I.
The evaluation shows that RPD-MSE outperforms all other variants, closely followed by RPD-L1. The distilled policy quickly surpasses Octo and OpenVLA to become a refined policy. RPD-MSE with OpenVLA performs slightly worse than the distillation with Octo. The PPO baseline also converges to similar success rates as RPD-MSE and RPD-L1 of around 80%, but learns substantially slower and shows much larger fluctuations in the training process. RPD-BC slightly surpasses Octo’s performance threshold but stays at that level and does not refine the policy. Lastly, PPD fails to lift the policy above the Octo baseline.
The reason why our PPD baseline does not manage to learn the task successfully is likely that it requires the teacher and student policy distributions to be of similar shape for the KL divergence to converge. However, we found that the action distributions of the VLA policies are sometimes bimodal, which results in non-optimal parameters for the approximated normal distributions that we use to compute the KL divergence in PPD. We also speculate that in our setup PPD depends strongly on the teacher’s performance. Lastly, PPD may also be sensitive to PPO hyperparameters, which we have not investigated. We did, however, modify the parameter $\lambda$, but observed no substantial change in the results.
@@
Figure 3: Average success rates during training for vanilla PPO, PPD, and three different RPD variants: RPD-MSE, RPD-L1, and RPD-BC, which all distill from Octo over five runs with different seeds. Shaded areas indicate standard deviations. We also evaluated RPD-MSE on OpenVLA. Due to computing constraints, we could only perform a single training run. The performance of the fine-tuned VLAs is indicated by dashed lines.
@@ IV-B RPD on Different Tasks
Figure 4: Success and reward curves for RPD with Octo and for vanilla PPO for both dense and sparse rewards on the six different ManiSkill3 tasks that are part of the fine-tuning dataset. The values are averaged over five runs with different seeds and recorded in an evaluation environment. Standard deviations are indicated by shaded areas. All training runs were performed with the same hyperparameters from the ManiSkill3 PPO baseline. For all tasks, RPD outperforms Octo quickly and converges faster than vanilla PPO. In some cases, it even finds good policies when vanilla PPO fails. This effect increases for sparse rewards.
Fig. 4 shows the training curves for RPD-MSE with Octo on six different manipulation tasks that were part of the training dataset. In all six cases, RPD learns faster than the PPO baseline and often converges to higher success rates. This increased performance can also be seen in the reward plots. For RollBall and StackCube, the PPO baseline fails to solve the tasks with the hyperparameters from ManiSkill3, whereas RPD achieves at least partial success with the same parameters. The decreased standard deviation also shows that RPD is generally more robust than the vanilla PPO baseline. These results demonstrate that RPD improves PPO’s training speed consistently and over many tasks while surpassing its final performance. It also refines the distilled policy on dense rewards.
Figure 5: Success and reward curves for RPD with Octo and for vanilla PPO on tasks that are not part of Octo’s fine-tuning dataset. The curves show average results for five training runs on different seeds and are recorded in evaluation environments. The shaded areas indicate standard deviations. Note that the Octo baseline does not represent the correct reward in the dense reward setting as it solves the tasks without tool use. See Sec. III for details.
@@ IV-C Sparse Rewards and Sample Efficiency
ManiSkill3 defines its sparse rewards as follows: 1-1- 1 for failed cases, 111 for success, and 0 otherwise. This means that non-zero reward is only given at the end of the episode, which is different from dense rewards where every action usually receives a non-zero reward that is shaped to guide the agent towards a certain behavior. Sparse reward problems are often a big challenge for RL algorithms, as relevant information about the agent’s performance is only obtained once the task has been solved successfully at least one time. Thus, in the beginning, the agent needs to complete an episode successfully just through random exploration, which makes the training process very inefficient compared to RL with dense rewards. As a result, it can occur that an agent learns a task perfectly fine with dense rewards but fails in the corresponding sparse reward setting, as can be seen for PPO in Fig. 4 for some of the tasks.
Fig. 4 also shows training runs with the six selected ManiSkill3 tasks for sparse rewards. Note that the resulting reward values cannot be directly compared to the dense reward setting. The only modification we applied to the hyperparameters from the ManiSkill3 baselines was to set $\gamma$ from 0.80.80.8 to 0.990.990.99, which avoids vanishing returns for early episode steps. In all six cases, RPD on sparse rewards consistently outperforms vanilla PPO. In most tasks, the baseline completely fails to learn the task, whereas RPD succeeds. It makes sense that RPD can realize its full potential on sparse rewards as it is even more important in this setting to have a teacher policy that enables the student to successfully complete episodes and, thus, retrieve environment rewards that drive the learning process. In summary, RPD drastically improves sample efficiency for on-policy RL in both sparse and dense reward cases.
@@ IV-D Generalization to New Environments
We also evaluated RPD with Octo on tasks that were held out of the VLA fine-tuning dataset to test its generalization capabilities. Ultimately, RPD is limited by the performance of the teacher and, in this case, by the cross-task generalization of the VLA. Still, it is worth investigating to what extent the degraded VLA performance will influence the RPD training. We study this question by looking at task variations, which means they are out-of-distribution with respect to the fine-tuning dataset. More specifically, we evaluate RPD’s performance on the hold-out tasks PullCubeTool and PokeCube, as well as on a variation of the PushCube task with a different camera perspective.
@@
VLAs have traditionally been very sensitive to camera adjustments, leading to severe performance degradation [9]. To counteract performance drops, VLAs often require fine-tuning with camera data from the new angle, camera model, etc. This problem also occurs with our fine-tuned versions of Octo and OpenVLA: When we only slightly change the camera angle in the PushCube task, as shown in Fig. 6, the success rates of OpenVLA drops from 27% to 4.5%. Octo even drops from 67% to 0%. The latter is likely a result of the less powerful vision encoder that is utilized in Octo that may overfit the training data.
Fig. 7 shows the success rates of vanilla PPO and RPD-MSE for both Octo and OpenVLA on the changed camera perspective from Fig. 6 on the PushCube task. PPO performs similarly on our task variation as in the original PushCube task with sparse rewards, see Fig. 4. This is expected because the agent’s visual encoder is learned from scratch in the RL training. Accordingly, there is no bias towards a specific camera angle as long as the provided information is sufficient to solve the task. Both RPD variants manage to outperform PPO.
@@
Figure 7: Average evaluation success rates of vanilla PPO and of RPD-MSE with both Octo and OpenVLA on a variation of the PushCube task with sparse rewards over five different seeds. Shaded areas indicate standard deviations. The camera perspective was changed as shown in Fig. 6 to investigate the distillation performance of the degraded VLAs.
@@ IV-E Limitations
The main limitation of RDP is the performance of the VLA on the task at hand. If the VLA cannot provide any meaningful actions due to domain shift or because the task was not included in the training dataset, RPD can be seen as a new exploration component in the loop. Depending on the resulting action distribution, the VLA may provide bias that is useful for learning the task or could improve the exploration of the environment. However, if the actions generated by the VLA cannot provide such helpful bias, RPD may also decrease the RL agent’s performance. As reported for the training with dense rewards on the hold-out tasks, if the task and no variation of it have ever been seen by the VLA, its actions hinder the PPO training in RPD, and vanilla PPO performs better. Similarly, when PPO is already tuned to a task or the task is very easy, PPO is already sample-efficient, and RPD may only provide limited improvement.
The method is platform-agnostic and can also be employed on a physical setup. However, despite RPD’s gains in sample efficiency, online RL still needs millions of steps to converge for more complex tasks, which makes it challenging on physical platforms. For example, sim-to-real methods could be employed to bridge this gap, as described below.
@@ V Conclusion and Future Work
This work introduced RPD, a novel on-policy RL-based approach for distilling task knowledge from generalist VLAs into refined expert policies. To the best of our knowledge, RPD is the first approach in this direction. Our experiments show that RPD not only makes the RL training more efficient by guiding its exploration through the actions of the VLA, but also that it can leverage task knowledge from the VLA even in situations where the generalist policy fails, such as when the camera perspective changes. Thus, RPD can help to improve the usability and accessibility of VLAs. As new VLAs become available [5], we expect the performance of RPD to improve even further.
Due to the high data requirements of RL, we evaluated RPD in simulation, which is why we fine-tuned two state-of-the-art generalist VLA policies, Octo and OpenVLA, on simulation datasets. As a side contribution, our work shows that VLAs can be successfully utilized in experiments in simulated environments. This provides a promising perspective for scientists who have no access to a physical robot setup to conduct research on VLAs. An interesting direction for future research is to combine RPD with sim-to-real methods [33] for deployment on real robots.
@@ Acknowledgement
This work is supported by the project GeniusRobot funded by the German Federal Ministry of Education and Research (BMBF grant no. 01IS24083). The authors acknowledge the HPC resources provided by the Erlangen National HPC Center (NHR@FAU) under the BayernKI project no. v108be.